# MGMCpaper.bib

### v1

@@ 5.2 Performance
shifted Laplace $\mathcal{A}^{(\mathrm{SL})}=-\Delta+\kappa^{2}I$ in $d=2$ dimensions, FEM discretisation
grid size IACT time / sample time / indep. sample
Gibbs MGMC Gibbs MGMC Chol. Gibbs MGMC
$32\times 32$ $3.1\pm 0.4$ $1.12\pm 0.12$ $0.15$ $0.22$ $0.08$ $0.46$ $0.25$
$64\times 64$ $10.5\pm 2.1$ $1.13\pm 0.12$ $0.62$ $0.91$ $0.39$ $6.47$ $1.04$
$128\times 128$ $47.3\pm 16.4$ $1.15\pm 0.13$ $2.53$ $3.71$ $1.79$ $119.93$ $4.28$
$256\times 256$ $95.1\pm 42.2$ $1.18\pm 0.14$ $10.18$ $15.17$ $9.14$ $968.49$ $17.96$
$512\times 512$ $264.2\pm 169.1$ $1.21\pm 0.15$ $43.03$ $66.69$ $47.23$ $11370.84$ $80.53$
shifted Laplace $\mathcal{A}^{(\mathrm{SL})}=-\Delta+\kappa^{2}I$ in $d=3$ dimensions, FD discretisation
grid size IACT time / sample time / indep. sample
Gibbs MGMC Gibbs MGMC Chol. Gibbs MGMC
$16\times 16\times 16$ $2.6\pm 0.3$ $1.32\pm 0.19$ $0.57$ $0.74$ $0.48$ $1.47$ $0.98$
$32\times 32\times 32$ $4.7\pm 0.7$ $1.20\pm 0.14$ $5.20$ $7.62$ $12.58$ $24.60$ $9.17$
$48\times 48\times 48$ $10.3\pm 2.0$ $1.26\pm 0.17$ $24.87$ $30.97$ $82.81$ $257.20$ $39.17$
$64\times 64\times 64$ $20.6\pm 5.3$ $1.28\pm 0.17$ $50.11$ $67.89$ $317.23$ $1030.82$ $87.00$
squared shifted Laplace $\mathcal{A}^{(\mathrm{SSL})}=(-\Delta+\kappa^{2}I)^{2}$ in $d=2$ dimensions, FD discretisation
grid size IACT time / sample time / indep. sample
Gibbs MGMC Gibbs MGMC Chol. Gibbs MGMC
$32\times 32$ $22.4\pm 5.9$ $2.22\pm 0.26$ $0.15$ $0.36$ $0.10$ $3.45$ $0.81$
$64\times 64$ $3401.7\pm 4757.4$ $3.35\pm 0.43$ $0.64$ $1.50$ $0.52$ $2167.88$ $5.02$
$128\times 128$ $1976.3\pm 2448.2$ $2.69\pm 0.35$ $2.61$ $6.15$ $2.49$ $5148.96$ $16.55$
$256\times 256$ $2573.9\pm 3415.3$ $3.23\pm 0.40$ $10.59$ $25.43$ $16.13$ $27265.49$ $82.01$
$512\times 512$ $1682.9\pm 1991.1$ $3.94\pm 0.57$ $45.09$ $110.39$ $81.68$ $75871.91$ $435.17$
@@ 6 Conclusion
In this paper we have presented a rigorous analysis of the Multigrid Monte Carlo approach originally proposed by Goodman and Sokal in [goodman1989multigrid]. We have extended the analysis to the important case of conditioned Gaussian random fields in a linear Bayesian setting. We have shown that the rate of convergence to the targest distribution and the integrated autocorrelation time is grid-independent. To achieve this, we relied on the inherent connection between solvers for sparse linear systems and samplers for multivariate normal distributions discussed in [Fox.C_Parker_2017_AcceleratedGibbsSampling]. This allowed us to bound convergence rates with tools from standard multigrid theory which, however, had to be extended to the Bayesian setup considered here. Our cost-analysis confirms that the cost for one MGMG update grows linearly in the number of unknowns. As a consequence, the algorithm is near-optimal in the sense that the cost for generating a single independent sample is also proportional to the number of unknowns.
The theoretical results are confirmed by a set of numerical experiments, which demonstrate that MGMC is robust and efficient as the resolution increases. On finer grids, MGMC is always significantly faster than a standard Gibbs sampler. In three dimensions and on larger lattices we were also able to beat the Cholesky method, which demonstrates that our approach is particularly promising in higher dimensions. We are able to sample very rough fields with precision operators $-\Delta+\kappa^{2}I$ and $(-\Delta+\kappa^{2}I)^{2}$ without problems, but find that MGMC is most efficient in the former case.
@@
There are several ways of extending the work in this paper. While we only considered relatively small problems here, significantly larger problems can be simulated with the parallel implementation which will be described in a subsequent publication. For simplicity, we have limited ourselves to a linear setting where in particular the quantity of interest and the observation operators are bounded linear functionals. It would be interesting to consider the more general case where the likelyhood is given by $\propto\exp\left[-\frac{1}{2}(\mathcal{B}_{\text{NL}}(\theta)-\mu)^{\top}\Gamma^{-1}(\mathcal{B}_{\text{NL}}(\theta)-\mu)\right]$ for some non-linear $\mathcal{B}_{\text{NL}}$; the resulting probability density will be log-concave. Similarly, the generalisation to non-Gaussian priors could be considered. While the original MGMC algorithm in [goodman1989multigrid] is written down for the general non-Gaussian case and should cover these cases in principle, in practice it will likely have to be adapted to be efficient.
@@ A.5 Convergence of distributions
Similarly, $\lim_{m\to\infty}{D_{\text{KL}}\bigl{(}(A_{L}^{-1}f_{L},A_{L}^{-1})\parallel(\text{$\mu_{m}$, $\Sigma_{m}$})\bigr{)}=0}$ holds, where we again note $\mathrm{det}(\Sigma_{m})\neq 0$ for sufficiently large $m$. In view of the Pinsker’s inequality, the total variation distance converges to 00 as well, which implies the weak convergence. The weak convergence can be checked directly with the convergence of characteristic function. [^1] Since $\theta_{L}^{(m)}$ is a normal random variable, the characteristic function $\varphi_{m}(x):=\varphi_{m,L}(x)$ of $\theta_{L}^{(m)}$ is given by $\varphi_{m}(x)=\exp(\mathrm{i}\mu_{m}^{\top}x-\frac{1}{2}x^{\top}\Sigma_{m}x)$ with $\mu_{m}:=\mathbb{E}[\theta_{L}^{(m)}]$ and $\Sigma_{m}:=\mathrm{Cov}(\theta_{L}^{(m)})$. The first two statements imply $\lim_{m\to\infty}\varphi_{m}(x)=\varphi(x):=\exp\Bigl{(}\mathrm{i}(A_{L}^{-1}f_{L})^{\top}x-\frac{1}{2}x^{\top}A_{L}^{-1}x\Bigr{)}\qquad\text{for any $x\in\mathrm{\mathbb{R}}$,}$ but we have $|\varphi_{m}(x)-\varphi(x)|$ $=\Bigl{|}\exp\Bigl{(}\mathrm{i}(\mu_{m}-A_{L}^{-1}f_{L})^{\top}x-\frac{1}{2}x^{\top}(\Sigma_{m}-A_{L}^{-1})x\Bigr{)}-1\Bigr{|}\leq|(\mu_{m}-A_{L}^{-1}f_{L})^{\top}x|+|\frac{1}{2}x^{\top}(\Sigma_{m}-A_{L}^{-1})x|$ and thus $\theta_{L}^{(m)}$ converges to a Gaussian random variable with mean $A_{L}^{-1}f_{L}$ and covariance $A_{L}^{-1}$ in distribution.
@@ A.9 Optimality of MGMC
From (A.49) and (A.50) we read off the constants $C_{1,1}:=C_{F}(C_{0}+C_{f})$, $C_{1,2}=C_{F}(1+C_{0})$ and $C_{2}=\log(C_{A}+\nu)-\log(C_{A})>0$. ∎

### v2

@@ 6 Conclusion
In this paper we have presented a rigorous analysis of the Multigrid Monte Carlo approach originally proposed by Goodman and Sokal in [goodman1989multigrid]. We have extended the analysis to the important case of conditioned Gaussian random fields in a linear Bayesian setting. We have shown that the rate of convergence to the targest distribution and the integrated autocorrelation time is grid-independent. To achieve this, we relied on the inherent connection between solvers for sparse linear systems and samplers for multivariate normal distributions discussed in [Fox.C_Parker_2017_AcceleratedGibbsSampling]. This allowed us to bound convergence rates with tools from standard multigrid theory which, however, had to be extended to the Bayesian setup considered here. Our cost-analysis confirms that the cost for one MGMG update grows linearly in the number of unknowns. As a consequence, the algorithm is optimal in the sense that the cost for generating a single independent sample is also proportional to the number of unknowns.
The theoretical results are confirmed by a set of numerical experiments, which demonstrate that MGMC is robust and efficient as the resolution increases. On finer grids, MGMC is always significantly faster than a standard Gibbs sampler. In three dimensions and on larger lattices, it beats even the Cholesky method, which demonstrates that our approach is particularly promising in higher dimensions. We are able to sample very rough fields with precision operators $-\Delta+\kappa^{2}I$ and $(-\Delta+\kappa^{2}I)^{2}$ without problems, but find that MGMC is most efficient in the former case.
@@
There are several ways of extending the work in this paper. While we only considered relatively small problems here, significantly larger problems can be simulated with the parallel implementation which will be described in a subsequent publication. For simplicity, we have also limited ourselves to a linear setting where in particular the quantity of interest and the observation operators are bounded linear functionals. It would be interesting to consider more general cases with likelihoods given by $\propto\exp\left[-\frac{1}{2}(\mathcal{B}_{\text{NL}}(\theta)-\mu)^{\top}\Gamma^{-1}(\mathcal{B}_{\text{NL}}(\theta)-\mu)\right]$ for some non-linear $\mathcal{B}_{\text{NL}}$. Similarly, the generalisation to non-Gaussian priors could be considered. While the original MGMC algorithm in [goodman1989multigrid] is written down for the general non-Gaussian case and these cases should be covered in principle, in practice it will likely have to be adapted to be efficient.
@@ A.5 Convergence of distributions
Similarly, $\lim_{m\to\infty}{D_{\text{KL}}\bigl{(}(A_{L}^{-1}f_{L},A_{L}^{-1})\parallel(\text{$\mu_{m}$, $\Sigma_{m}$})\bigr{)}=0}$ holds, where we again note $\mathrm{det}(\Sigma_{m})\neq 0$ for sufficiently large $m$. In view of the Pinsker’s inequality, the total variation distance converges to 00 as well, which implies the weak convergence. The weak convergence can be checked directly with the convergence of characteristic function. ∎
@@ A.9 Optimality of MGMC
From (A.49) and (A.50) we read off the constants $C_{1,1}:=C_{F}(C_{0}+C_{f})$, $C_{1,2}=C_{F}(1+C_{0})$ and $C_{2}=\log(C_{A}+\nu)-\log(C_{A})>0$.
Finally, noting that $u:=\widetilde{A}^{-1}f$ satisfies the problem $a(u,v)+b(u,v)=\langle f,v\rangle_{H}$ for all $v\in H$, Assumption 2.2 implies that $P_{L}u_{L}:=P_{L}\widetilde{A}_{L}^{-1}f_{L}$ with $f_{L}=(\langle f,\phi_{j}^{L}\rangle_{H})_{j=1,\dots,n_{L}}$ converges to $u$ in $V$ as $L\to\infty$. Hence,
$\lim_{L\to\infty}\mathbb{E}[F_{L}^{\top}\theta_{L}]=\lim_{L\to\infty}\mathcal{F}P_{L}\widetilde{A}_{L}^{-1}f_{L}=\mathcal{F}\widetilde{A}^{-1}f$
Similarly, $\lim_{L\to\infty}\mathrm{Cov}(F_{L}^{\top}\theta_{L})=\lim_{L\to\infty}\mathcal{F}P_{L}\widetilde{A}_{L}^{-1}F_{L}=\lim_{L\to\infty}\mathcal{F}\widetilde{A}^{-1}\mathcal{F}$, where we used $F_{L}=(\langle\mathcal{F},\phi_{j}^{L}\rangle_{H})_{j=1,\dots,n_{L}}$. Thus, the characteristic function of $(F_{L}^{\top}\theta_{L}^{*})_{L}$ converges to that of $(\widetilde{v},\mathcal{F})$ pointwise. Now the proof is complete. ∎